ESTA Inc. has a beta of 2, the market value of equity of $50M and the market value of assets of $83M. The corporate tax rate is 20%. The expected risk premium on stock market is 6.3%, while the current interest rate on T-bonds is 4%. ESTA pays an 8% coupon rate on its 30-year bonds. Bonds make semiannual payments. The current price of these bonds is $1080. Estimate ESTA cost of capital.
Answer #1

Assets=83
Equity=50
Debt=Assets-Equity=83-50=33

Cost of equity=risk free rate+beta*market risk premium=4%+2*6.3%=16.6%

Semi annual yield to maturity:
Using financial calculator
N=30*2=60
PMT=8%*1000/2=40
PV=-1080
FV=1000
CPT I/Y=3.6683%

Tax rate=20%

Cost of capital=33/83*3.6683%*2*(1-20%)+50/83*16.6%=12.3336%